kekeman:

Find the product of (x + (3+5i))^2

Rylee88:

I don't know exactly but, x and I are your variables BOTH. (3+5i) Find the variable first then do 3+5_=__ after times it by 2 square. Then do x+ your answer. I think anyways. Hope this helps, If you need more help let me know and I'll try my best.

Vocaloid:

In the context of this problem, i is not a variable. i is a constant. remember binomial expansion: (x+a)^2 = x^2 + 2ax + a^2 in this case, treat the entire (3+5i) as your a expression

surjithayer:

\[(x+(3+5\iota))^2\] \[=[(x+3)+5\iota]^2\] \[=(x+3)^2+(5 \iota)^2+2(x+3)(5\iota)\] =?

kekeman:

(x + (3 + 5i))^2 (x + 3 + 5i)(x + 3 + 5i) x^2 + 3x + 5xi + 3x + 9 + 15i + 5xi + 15i + 25i^2 x^2 + 6x + 10xi + 9 + 30i +25i^2 x^2 + 6x + 10xi - 16 + 30i

Vocaloid:

that's correct

kekeman:

Oh fr thanks